OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3(x – 5); (–4, 3)

OpenStudy (anonymous):

m2=-3,y-3=-3(x+4)_________________>y-3=-3x-12____>y=-3x-9

OpenStudy (anonymous):

Thanks. Couldn't figure this one out.